OpenStudy (anonymous):

-4 square root of -54 using i notationa as needed

OpenStudy (anonymous):

\[-4\sqrt{-54}\]

OpenStudy (anonymous):

\[\LARGE -4\sqrt{-54}=-4i\cdot\sqrt{ 54}=-4i\cdot 3\sqrt6 =??\]

OpenStudy (anonymous):

\[-5\sqrt{-24}\]

OpenStudy (anonymous):

how about that one, sorry i wrote the wrong #s

OpenStudy (anonymous):

follow the steps I did in post above.